OpenStudy (anonymous):

Solve using elimination method: 0.3x - 0.2y = 4 0.4x + 0.5y = 1

OpenStudy (anonymous):

My answer has to be in an ordered pair and can be used by integers or fractions. It also can have infinitely many solutions or none.

OpenStudy (amistre64):

elimination is such a cruel method...

OpenStudy (anonymous):

I do not understand elimination method at all. I can do substitution but not this stuff.

OpenStudy (anonymous):

i got (7.14,9.286) it's been awhile since ive done this

OpenStudy (amistre64):

you simply eliminate a variable by multiplying or modifying one equation to be the opposite of the other... 0.3x - 0.2y = 4 0.4x + 0.5y = 1

OpenStudy (anonymous):

Mjrizon, that is not correct.

OpenStudy (amistre64):

3/10x - 2/10y = 4 4/10x + 5/10y = 1 (-3/4) 3/10x - 2/10y = 4 -3/10x -15/40y = -3/4

OpenStudy (amistre64):

3/10x - 8/40y = 16/4 -3/10x -15/40y = -3/4 ---------------------- -23/40y = 13/4

OpenStudy (anonymous):

oh ok sorry. forgot to multiply the right side of the equations by the multipliers.

OpenStudy (anonymous):

please dont do all her homework.

OpenStudy (anonymous):

hi amistre, i figured out the surface area problem

OpenStudy (amistre64):

-23/40y = 13/4 y = 13/4 -40/23 y = 13/1 -10/23 = -130/23

OpenStudy (anonymous):

I am stuck on three. Cant people help me with them.

OpenStudy (amistre64):

thats great :)

OpenStudy (anonymous):

blondie, you are being graded or not , be honest

OpenStudy (anonymous):

oh ok sorry. forgot to multiply the right side of the equations by the multipliers.

OpenStudy (anonymous):

yes i will be graded this is a checkpoint. i have the test for this next week though so it will be all new problems.

OpenStudy (anonymous):

so youre cheating pretty much

OpenStudy (amistre64):

is my "y" value good?

OpenStudy (anonymous):

amistre, give her hints :) not answers

OpenStudy (anonymous):

No im asking for help on these last three problems.

OpenStudy (anonymous):

no you didnt ask for help. i read your log

OpenStudy (anonymous):

you wanted the answer

OpenStudy (anonymous):

youre not doing any work

OpenStudy (anonymous):

Okay no i didnt just do any homework whatever. Ill go elsewhere to get help from people then because I dont cheat on my homework

OpenStudy (amistre64):

4x 5 (-130) -- + -------- =1 10 10(23)

OpenStudy (anonymous):

i dont mean to be rude but i dont see any work coming from you

OpenStudy (anonymous):

If i am not doing any of my work then how is my grade an 87.5 for this checkpoint right now?

OpenStudy (anonymous):

amistre, dont kill yourself

OpenStudy (anonymous):

blondie ok . how did you get those answers

OpenStudy (anonymous):

How did i get the other answers for my checkpoint? I was able to figure them out because it wasnt elimination method. It was substitution and graphing.

OpenStudy (anonymous):

youre not asking questions, youre just being spoon fed answers . you would know about that

OpenStudy (amistre64):

4x -65 -- + -------- =1 10 (23) 4x 23 + 65 -- = --------- 10 23 22 10 220 x = ------ = ---- if i didnt mess up along the way 23 23

OpenStudy (anonymous):

this is not the place to post a thousand annoying homework problems,

OpenStudy (anonymous):

post interesting problems

OpenStudy (anonymous):

Yep sure would since my grades are at ah lets see an 89% overall. and I ASKED 3 NOT A THOUSAND

OpenStudy (anonymous):

I think someone should mind their own business.

OpenStudy (anonymous):

3 too many. 1 is a lot

OpenStudy (anonymous):

well im looking out for my buddies

OpenStudy (anonymous):

50 percent of the questions dont get answered, what makes you think youre entitled to it

OpenStudy (anonymous):

90 percent of MY QUESTIONS HAVE NOT BEEN ANSWERED

OpenStudy (anonymous):

anyways, amistre , just word to the wise

OpenStudy (amistre64):

(3/10)(220/23) - (2/10)(-130/23) = 4 (3/1)(22/23) - (2/1)(-13/23) = 4 (66/23) + (26/23) = 4 92/23 = 4 ....4(23) = 92 that ones good

OpenStudy (amistre64):

word to the wise? looks left...looks right.... :)

OpenStudy (anonymous):

should i check your guys work? hmmm, i wonder if you can do this with determinants quickly

OpenStudy (amistre64):

ess one question...and i most likely got it wrong, but might have gotten it right ;)

OpenStudy (anonymous):

I agree with cantorset.

OpenStudy (anonymous):

Although unfortunately people do NOT use this website to post interesting problems, although I guess it is not technically meant for that. If only it were :( So much better.

OpenStudy (anonymous):

i mean youre free to do anything you want. i was just saying, dont exhaust yourself doing solutions without feedback

OpenStudy (anonymous):

newton, yeah, ideally

OpenStudy (amistre64):

oh i agree cantor.... i agree :)

OpenStudy (anonymous):

oh believe me i wont come back to this website. I will find help elsewhere as i said. You can be a wingspan to someone else!!!

OpenStudy (amistre64):

i was wondering if I could do the surface area of that torus with polar stuff....

OpenStudy (anonymous):

i guess interesting is relative/subjective though

OpenStudy (anonymous):

blondie, thats fine

OpenStudy (anonymous):

You were a little harsh to her, specifically, though xD Though I do see that kind of thing on here a lot.

OpenStudy (anonymous):

amistre, yes you can. i want to make a pdf of all the different ways to find surface area

OpenStudy (amistre64):

what i came up with looked like a nightmare to integrate...but it was fun trying lol

OpenStudy (anonymous):

i apologize im harsh today , but i do help

OpenStudy (anonymous):

amistre oh its a fun problem, i will come back to it im going to help avnis , one sec

OpenStudy (amistre64):

wish i could stay and play, but its 3am here and i gotta be able to get up and get my kids in a few hours;) nite nite yall ..

OpenStudy (anonymous):

ok will post tomorrow im happy i figured it out on my own. nobody helped me here :(

OpenStudy (amistre64):

your smart enough to do it on your own ;)

OpenStudy (anonymous):

i peaked your interest , thats good

OpenStudy (anonymous):

well i was stumped. but then i figured out it has to be integral 2pi x dS, to revolve about y axis integral 2pi y dS revolves about x axis ( and that makes a sphere)

OpenStudy (anonymous):

thats why we got 4pi*r^2 something